OpenStudy (anonymous):

give any 2 example sums for each of the following: (a+b)whole square (a-b)whole square (a+b)(c+d) (a-b)(c+d) (a-b)(c-d)

OpenStudy (aravindg):

Take any 2 numbers and pug in eg: (2+1)^2=2^2+1^2+2*2*1 9=4+1+4=9

OpenStudy (aravindg):

What do u mean sums for 2 sums?

OpenStudy (aravindg):

You mean examples?

OpenStudy (aravindg):

It's too simple. Just take any two numbers and apply the expansion.

OpenStudy (aravindg):

(1+root(3)}^2 (4+7)(4-7) Just make an attempt. It isn't hard.

OpenStudy (aravindg):

Good :)

OpenStudy (anonymous):

\[(2+9 )^{2}\]

OpenStudy (anonymous):

\[(36-28)^{2}\]

OpenStudy (anonymous):

i want a sum for every formula, which means formula used for the sum

OpenStudy (anonymous):

\[( 6 +8)(45+66)\]

OpenStudy (anonymous):

in the sense i want questions

OpenStudy (anonymous):

is those ok?

OpenStudy (anonymous):

i meant per formula 2 questions

OpenStudy (anonymous):

(2587+2548) (5986-1111)

OpenStudy (anonymous):

hey guys i will giv a meda, but please help

OpenStudy (aravindg):

This question is over. You should not ask for direct answers.

OpenStudy (anonymous):

i gave aravind a medel for no reason lol

OpenStudy (aravindg):

Haha like a boss >.> @hey,im,miranda

OpenStudy (anonymous):

;D

OpenStudy (aravindg):

You got your help. Now u are asking us to do the entire work. It doesn't work that way.

OpenStudy (anonymous):

:)